A system of linear equations has the augmented matrix


"\left(\begin{array}{cccc}1&1&k&:&1\\-1&1&2k&:&-3\\k&1&1&: &k+1\end{array}\right)"

where k

k is a constant. What is the value of k such that the system above has no solution?


  1. "k=-1"
  2. "k=-2"
  3. "k=0"
  4. "k=1"
1
Expert's answer
2021-08-30T19:07:46-0400

Solution:

For no solution, "\det \begin{pmatrix}1&1&k\\ \:-1&1&2k\\ \:k&1&1\end{pmatrix}=0"

"\Rightarrow 0=1\cdot \det \begin{pmatrix}1&2k\\ 1&1\end{pmatrix}-1\cdot \det \begin{pmatrix}-1&2k\\ k&1\end{pmatrix}+k\det \begin{pmatrix}-1&1\\ k&1\end{pmatrix}"

"\Rightarrow 0=1\cdot \left(1-2k\right)-1\cdot \left(-1-2k^2\right)+k\left(-1-k\right)
\\\Rightarrow 0=k^2-3k+2
\\\Rightarrow k^2-2k-k+2=0
\\\Rightarrow (k-2)(k-1)=0
\\\Rightarrow k=2,k=1"

Thus, option 4 is correct.